OpenStudy (anonymous):

How do I convert r=2-cos(3theta) to a cartesian equation?

OpenStudy (anonymous):

@agent0smith

OpenStudy (anonymous):

Also, the question was originally given in parametric form of (r,theta)=(2-cos(t3pi),tpi) Did I eliminate the parameters correctly?

OpenStudy (anonymous):

@sourwing

ganeshie8 (ganeshie8):

looks good, multiply "r" both sides

ganeshie8 (ganeshie8):

and use the trig identiy : cos(3x) = 4cos^3x - 3cosx

ganeshie8 (ganeshie8):

\( r=2-\cos(3\theta) \) \(r^2 = 2r - r[4\cos^3\theta - 3\cos \theta]\) \(x^2+y^2 = 2\sqrt{x^2+y^2} - r[4\frac{x^3}{r^3} - 3\frac{x}{r}]\) \(x^2+y^2 = 2\sqrt{x^2+y^2} - 4\frac{x^3}{x^2+y^2} - 3x\)

OpenStudy (anonymous):

Wow, that looks complicated. You're sure I got rid of the parameters correctly?

ganeshie8 (ganeshie8):

looks okay to me good to have checked it wid multiple eyes @agent0smith

OpenStudy (anonymous):

Ok, well thank you so much!

OpenStudy (agent0smith):

\[\Large r=2- 4\cos^3 \theta - 3\cos \theta\] there's no simple way to do it... multiply both sides by r^3 and i'm not gonna explain all steps \[\Large (r^2)^2=2r^2r- 4 (r\cos \theta)^3 - 3r^2 r\cos \theta\] \[\Large (x^2 + y^2)^2=2 (x^2+y^2) \sqrt{ x^2+y^2 } - 4 x^3 - 3x(x^2+y^2) \]

OpenStudy (agent0smith):

\[\Large (x^2 + y^2)^2=2 (x^2+y^2)^{3/2} - 4 x^3 - 3x(x^2+y^2) \]

OpenStudy (anonymous):

Is there anyway to arrive at an equation in cartesian form just using the given parametric equation?

OpenStudy (agent0smith):

Maybe. Idk tho, it's weird lookin http://www.wolframalpha.com/input/?i=convert+to+rectangular+r%3D2-cos%283theta%29+

OpenStudy (anonymous):

Yup certainly is. Well thanks everyone for your help.

OpenStudy (agent0smith):

And based on how it looks in rectangular form, either mine or @ganeshie8's i doubt there's a simple representation (hence the usefulness of polar or parametric form)

OpenStudy (agent0smith):

Oh and @ganeshie8's is the same as mine if you multiply both sides by (x^2+y^2)